Five people decide to share a yacht. By bringing in an additional partner, they can lower the cost to each person by 7000 dollars. What is the total cost of the yacht?
Answer by ewatrrr(24785)


Hi
Let x represent the cost of the yacht
Question states*** this is a perfect ex of finding the words that are 'Set-Up' words
x/5 = x/6 + $7000
Solving for x
1/30x = $7000
x =$210,000
CHECKING our Answer***
$210,000/5 = $42,000
$210,000/6 = $35,000